OpenStudy (anonymous):

A triangle is made by cutting a diagonal on rectangle 33 1/3% longer than its width. If the area of the triangle is 54cm^2, determine the lengyh and width of the original rectangle *please include solution too.

OpenStudy (dumbcow):

33 1/3 % = 1/3 so if width of rectangle is w, then diagonal is (4/3)w Are of triangle is half the Area of rectangle --> w*h = 108 then use pythagorean thm w^2 + h^2 = (4/3)^2 w^2 h^2 = 7/9 w^2 h = (sqrt7 /3)w --> (sqrt7/3)w^2 = 108 --> w = 18/7 --> h = 6sqrt7/7

OpenStudy (dumbcow):



OpenStudy (dumbcow):

umm sorry , my final answers are wrong but all the steps are correct trying to solve in my head :|